# On an entropy inequality for quadratic forms and applications

### v1

@@ 2.1 Notations
Let $(\Omega,\mathfrak{B},P)$ be a probability space. A random variable $X$ is a measurable function from $(\Omega,\mathfrak{B},P)$ to another measuable space $(\Omega^{\prime},\mathfrak{B}^{\prime})$, on which a pushforward probability measure $\mu$ is induced, i.e., $\mu(B^{\prime})=P(X\in B^{\prime})$ for any $B^{\prime}\in\mathfrak{B}^{\prime}$. We will say that $\mu$ is the associated distribution measure of $X$. We denote the range of $X$ by $range(X)$, which collects all the possible values that $X$ can take. In this paper, all the measurable spaces will be taken to be subspaces of ${\mathbb{R}}^{m}$, equipped with the Borel $\sigma$-algebra or its sub-$\sigma$-algebra, and the Euclidean norm ||$\cdot$|| ⋅ |. The dimension $m$ is always regarded as an absolute constant. The range of any random variable is assumed to be a bounded set. When two random variables $X$ and $Y$ have the same distribution, we write $X\equiv Y$.
@@
For a measure $\mu$ and a Borel set $E$ with $\mu(E)>0$, we define the conditional measure of $\mu$ with respect to $E$ by $\mu_{E}(B)=\frac{\mu(B\cap E)}{\mu(E)}$ for any Borel set $B$. In particular, when $\mu=P$, the conditional probability measure given an event $E$ is denoted by $P_{E}$. For a random varaible $X$ and such an $E$, we denote by $X_{E}$ the random variable satisfying $X_{E}\equiv(X|E)$ (see Lemma B.2). That is to say,
@@
In this paper, by $f=O_{\kappa}(g)$, $f\ll_{\kappa}g$ or $g\gg_{\kappa}f$, we mean that there exists an absolute positive constant $C=C(\kappa)$ depending on the parameter(s) $\kappa$ such that $|f|\leq Cg$. All logarithms we use here have base $2$.
@@ 2.2 Entropy and basic properties
In above definition, the convention $00\log 0=$ 0 is used for convenience. We also say $\mu(B)H(\mu_{B},\mathcal{A})=0$ when $\mu(B)=0$, although $\mu_{B}$ is not defined in this situation. Intuitively, the entropy of a random variable will tell us how hard it is to determine a value. Before stating the first lemma, we introduce the following notation:
@@
Moreover, the function $\iota(x)=-x\log x$ is non-negative valued, bounded and concave on $[0,1]$. It is increasing on $[0,e^{-1}]$, and decreasing on $[e^{-1},1]$. With this in tow, we record some properties of entropy for partitions, whose proofs are simple exercises.
@@
(i) $H(\mu,\mathcal{A})\geq 0$ with equality holding if and only if $|Z(\mathcal{A})|=1.$
(ii) $H(\mu,\mathcal{A})\leq\log|Z(\mathcal{A})|$ when $Z(\mathcal{A})$ is a finite set, with equality holding if and only if $\mu(A)=1/|Z(\mathcal{A})|$ for all $A\in Z(\mathcal{A}).$
@@
(i) $H(X)\geq 0$ with equality holding if and only if $|range_{+}(X)|=1$.
(ii) $H(X)\leq\log|range_{+}(X)|$ when $range_{+}(X)$ is a finite set, with equality holding if and only if $X$ is uniformly distributed on $range_{+}(X)$.
@@
The following is known as the Plünnecke-Ruzsa inequality. A proof can be found several places, for example, [7, Theorem 3.11].
Lemma 2.9 (Plünnecke-Ruzsa inequality).
Let $X,Y_{1},Y_{2},\ldots,Y_{m}$ be independent discrete random variables. Then
$H(X+Y_{1}+\ldots+Y_{m})-H(X)\leq\sum\limits_{i=1}^{m}\big{(}H(X+Y_{i})-H(X)\big{)},$
$H(Y_{1}+\ldots+Y_{m})-H(X)\leq\sum\limits_{i=1}^{m}\big{(}H(X+Y_{i})-H(X)\big{)}.$
@@
Let $k\geq 1$ and $20<\lambda\leq 1/$ 2. Suppose that $p_{1},p_{2},\ldots,p_{k}$ are non-negative numbers satisfying $\sum\nolimits_{i=1}^{k}p_{i}\leq\lambda$. Then
$-\sum\limits_{i=1}^{k}p_{i}\log p_{i}\leq\lambda\log k+O(1).$
@@
which is attainable in the feasible region of $(p_{1},\ldots,p_{k+1})$. For any $11\leq j<i\leq k+$ 1, denote
@@
That is to say, when the other $k-1$ variables are given, the small $p_{i}$ is, the larger value $\tilde{f}$ and $f$ take. Now one can verify that the maximum value of $f$ is taken at $(p_{1},\ldots,p_{k+1})$ with
@@
$-\sum\limits_{i=1}^{k+1}p_{i}\log p_{i}\leq k\cdot\left(-\frac{\lambda}{k}\log\frac{\lambda}{k}\right)-(1-\lambda)\log(1-\lambda)\leq\lambda\log k+1.$
@@
$-\sum\limits_{i=1}^{k}p_{i}\log p_{i}\leq\lambda\log k+1+p_{k+1}\log p_{k+1}=\lambda\log k+O(1).$
@@ 2.3 Discretized entropy and properties
Combining Lemma B.2, one can verify that the statement in Lemma 2.5 also holds for $H_{n}$: If $X$ is a random variable and $\mathcal{B}$ is a finite partition. Then
@@
Let $\alpha,c_{1},c_{2}$ be real numbers with $\alpha\neq 0$ and $0<c_{2}<c_{1}$. Let $f(x)=\alpha x^{2}$ with domain $[-c_{1},-c_{2}]\cup[c_{2},c_{1}]$. Note that, by the mean value theorem,
@@
[…]
@@
Let $l\geq 2$ and $X_{1},X_{2},\ldots,X_{l}$ be random variables. Then for any $n\in\mathbb{N}$,
$H_{n}(X_{1}+X_{2}+\ldots+X_{l})=H(D_{n}(X_{l})+D_{n}(X_{2})+\ldots+D_{n}(X_{l}))+O(\log l).$
@@
Note that $\sum\nolimits_{i=1}^{l}D_{n}(X_{i})$ takes values in the dyadic rationals. One has
$H\left(\sum\limits_{i=1}^{l}D_{n}(X_{i})\right)=H\left(D_{n}\Big{(}\sum\limits_{i=1}^{l}D_{n}(X_{i})\Big{)}\right)=H_{n}\left(\sum\limits_{i=1}^{l}D_{n}(X_{i})\right).$
The conclusion then follows from Lemma 2.19 with the $O(l)$-Lipschitz function (see Example 2.12) given by
$(x_{1},\ldots,x_{l})\mapsto x_{1}+\ldots+x_{l}.$
Now we show a discretized version of the Plünnecke-Ruzsa inequality.
@@
Let $l\geq 1$ and $X,Y_{1},Y_{2},\ldots,Y_{l}$ be independent random variables. Then
$H_{n}(X+Y_{1}+Y_{2}+\ldots+Y_{l})-H_{n}(X)\leq\sum\limits_{i=1}^{l}\big{(}H_{n}(X+Y_{i})-H_{n}(X)\big{)}+O(\log l),$
and
$H_{n}(Y_{1}+Y_{2}+\ldots+Y_{l})-H_{n}(X)\leq\sum\limits_{i=1}^{l}\big{(}H_{n}(X+Y_{i})-H_{n}(X)\big{)}+O(\log l).$
Proof.
Note that $D_{n}(X),D_{n}(Y_{1}),\ldots D_{n}(Y_{l})$ are also independent. The conclusion follows from Lemmas 2.9 and 2.20. ∎
Corollary 2.22.
Let $X,Y,$ and $Z$ be independent random variables. Then, for all $n\in\mathbb{N}$,
$H_{n}(X)$ $\leq H_{n}(X+Y)+O(1),$
@@
We only show the proof of the third inequality involving $H_{n}(X-Y)$ here. The others are direct consequences of Lemma 2.21.
@@
By the independence of $X,Y,$ and $Z,$ (specifically, by invoking (8) and (9)) it follows that
$H_{n}(X-Y)+H_{n}(X,Y)+H_{n}(Z)\leq H_{n}(X+Z)+H_{n}(Y+Z)+H_{n}(X,Y)+O(1).$
The conclusion follows. ∎
@@
Let $X$ and $Y$ be independent real random variables and $k,l\in\mathbb{Q}$. There exist non-negative integers $d_{1}=d_{1}(k,l)$ and $d_{2}=d_{2}(k,l)$ such that
$H_{n}(kX+lY)\leq(d_{1}+d_{2}+1)H_{n}(X+Y)-d_{1}H_{n}(X)-d_{2}H_{n}(Y)+O_{k,l}(1).$
Here by $kX$ we mean the product of $k$ and the random variable $X$.
@@
by Corollary 2.22. Similar conclusion holds when $k=0$. In the following, we assume that $k,l\neq 0$.
@@
Indeed, when $k=l=1$, one can deduce (12) from the fact that
$H_{n}(X+Y)\geq\max\{H_{n}(X),H_{n}(Y)\}+O(1).$
When $k=1$ and $l=-1$, we take $X,Y,Z$ to be independent random variables such that $X$ and $Z$ are identically distributed. Then Corollary 2.22 gives
@@
Then (12) follows. Since $H_{n}(W)=H_{n}(-W)$ for any random variable $W$, the cases that $k=-1$ and $l=\pm 1$ are also satisfied.
@@
where the random variable $Z$ is taken such that $X,Y,Z$ are independent and $X$, $Z$ are identically distributed. Combining Corollary 2.22, we obtain that
@@
Combining (13) and the inductive hypothesis on $k$, one sees that the conclusion holds with some $d_{1}(k,l)$ and $d_{2}(k,l)$ satisfying
@@
By symmetry, we also have the induction on $l$, and the proof is completed for the situation that $k,l\in\mathbb{Z}^{\ast}$.
@@
for some $d_{1}=d_{1}(k^{\prime},l^{\prime})$ and $d_{2}=d_{2}(k^{\prime},l^{\prime})$. The lemma then follows. ∎
@@
for any $n\in\mathbb{N}$ and any i.i.d. random variables $X$, $Y$ with bounded range. However, when $k,l$ are $\mathbb{Q}$-linearly independent, such statement may fail. In Example A.6, we would show the following: For any $K>0$, there exists some $N\in\mathbb{N}$ such that: for any $n>N$, there is a random variable $X$ with range contained in $[0,1]$ satisfying
@@ 2.4 Frostman measures
Now we introduce Frostman measures and their basic properties.
Definition 2.24.
Let $s,C>0$. We say a Borel measure $\mu$ on $\mathbb{R}^{m}$ is $(s,C)$-Frostman if for any $x\in\mathbb{R}^{m}$ and $r>0$
@@
Let $X$ and $Y$ be independent real random variables that are both $(s,C)$-Frostman, where $0<s\leq 1$ and $C>0$. Then the following statements hold for $a,b,c,d\in\mathbb{R}$.
(i) The random variable $aX+bY$ is $(s,O_{a,b}(C))$-Frostman.
(ii) For any invertible linear change of variables $(x_{1},y_{1})=(ax+by,cx+dy)$, the random variable $(X_{1},Y_{1})=(aX+bY,cX+dY)$ is $(2s,O_{a,b,c,d}(C^{2}))$-Frostman.

### v2

@@ 2.1 Notations
Let $(\Omega,\mathfrak{B},P)$ be a probability space. A random variable $X$ is a measurable function from $(\Omega,\mathfrak{B},P)$ to another measuable space $(\Omega^{\prime},\mathfrak{B}^{\prime})$, on which a pushforward probability measure $\mu$ is induced, i.e., $\mu(B^{\prime})=P(X\in B^{\prime})$ for any $B^{\prime}\in\mathfrak{B}^{\prime}$. We will say that $\mu$ is the associated distribution measure of $X$. We denote the range of $X$ by $range(X)$, which collects all the possible values that $X$ can take. In this paper, all the measurable spaces will be taken to be subspaces of ${\mathbb{R}}^{m}$, equipped with the Borel $\sigma$-algebra or its sub-$\sigma$-algebra, and the Euclidean norm $|\cdot|$. The dimension $m$ is always regarded as an absolute constant. The range of any random variable is assumed to be a bounded set. When two random variables $X$ and $Y$ have the same distribution, we write $X\equiv Y$.
@@
For a measure $\mu$ and a Borel set $E$ with $\mu(E)>0$, we define the conditional measure of $\mu$ with respect to $E$ by $\mu_{E}(B)=\frac{\mu(B\cap E)}{\mu(E)}$ for any Borel set $B$. In particular, when $\mu=P$, the conditional probability measure given an event $E$ is denoted by $P_{E}$. For a random varaible $X$ and such an $E$, we denote by $X_{E}$ the random variable satisfying $X_{E}\equiv(X|E)$ (see Lemma A.2). That is to say,
@@
In this paper, by $f=O_{\kappa}(g)$, $f\ll_{\kappa}g$ or $g\gg_{\kappa}f$, we mean that there exists an absolute positive constant $C=C(\kappa)$ depending on the parameter(s) $\kappa$ such that $|f|\leq Cg$.
All logarithms we use here have base $2$. Moreover, we denote $\log_{+}x:=\max\{\log x,\,0\}$ for any $x>0$.
@@ 2.2 Entropy and basic properties
In above definition, the convention $0\log 0=0$ is used for convenience. We also say $\mu(B)H(\mu_{B},\mathcal{A})=0$ when $\mu(B)=0$, although $\mu_{B}$ is not defined in this situation. Intuitively, the entropy of a random variable will tell us how hard it is to determine a value. Before stating the first lemma, we introduce the following notation:
@@
Moreover, the function $\iota(x)=-x\log x$ is non-negative valued, bounded and concave on $[0,1]$. It is increasing on $0,e^{-1}]$, and decreasing on $e^{-1},1]$. With this in tow, we record some properties of entropy for partitions, whose proofs are simple exercises.
@@
(i) $H(\mu,\mathcal{A})\geq 0$ with equality holding if and only if $Z(\mathcal{A})=1.$
(ii) $H(\mu,\mathcal{A})\leq\log\#Z(\mathcal{A})$ when $Z(\mathcal{A})$ is a finite set, with equality holding if and only if $\mu(A)=1/\#Z(\mathcal{A})$ for all $A\in Z(\mathcal{A}).$
@@
(i) $H(X)\geq 0$ with equality holding if and only if $range_{+}(X)=1$.
(ii) $H(X)\leq\log\#\,range_{+}(X)$ when $range_{+}(X)$ is a finite set, with equality holding if and only if $X$ is uniformly distributed on $range_{+}(X)$.
@@
Let $k\geq 1$ and $0<\lambda\leq 1/2$. Suppose that $p_{1},p_{2},\ldots,p_{k}$ are non-negative numbers satisfying $\sum\nolimits_{i=1}^{k}p_{i}\leq\lambda$. Then
$\sum\limits_{i=1}^{k}p_{i}\log p_{i}\leq\lambda\log k+O(1).$
@@
which is attainable in the feasible region of $(p_{1},\ldots,p_{k+1})$. For any $1\leq j<i\leq k+1$, denote
@@
That is to say, when the other k1k-1 variables are given, the small $p_{i}$ is, the larger value $\tilde{f}$ and $f$ take. Now one can verify that the maximum value of $f$ is taken at $(p_{1},\ldots,p_{k+1})$ with
@@
$\sum\limits_{i=1}^{k+1}p_{i}\log p_{i}\leq k\cdot\left(-\frac{\lambda}{k}\log\frac{\lambda}{k}\right)-(1-\lambda)\log(1-\lambda)\leq\lambda\log k+1.$
@@
$\sum\limits_{i=1}^{k}p_{i}\log p_{i}\leq\lambda\log k+1+p_{k+1}\log p_{k+1}=\lambda\log k+O(1).$
@@ 2.3 Discretized entropy and properties
Combining Lemma A.2, one can verify that the statement in Lemma 2.5 also holds for $H_{n}$: If $X$ is a random variable and $\mathcal{B}$ is a finite partition. Then
@@
Let $\alpha,c_{1},c_{2}$ be real numbers with $\alpha\neq 0$ and $0<c_{2}<c_{1}$. Let $f(x)=\alpha x^{2}$ with domain $c_{1},-c_{2}]\cup[c_{2},c_{1}]$. Note that, by the mean value theorem,
@@
whenever $x,y\in[-c_{1},-c_{2}]$, or $x,y\in[c_{2},c_{1}]$. One asserts that $f$ is weakly $\max\left\{(2|\alpha|c_{2})^{-1},\,2|\alpha|c_{1}\right\}$-bi-Lipschitz.
@@
The following lemma says that the $n$-discretized operation makes no harm when we study entropy of sums of random variables.
@@
Let $k,l\geq 0$ and $X_{1},\ldots,X_{k}$, $Y_{1},\ldots,Y_{l}$ be real random variables. Then for any $n\in\mathbb{N}$,
$H_{n}(X_{1}+\ldots+X_{k},\,Y_{1}+\ldots,Y_{l})$
$\quad=H(D_{n}(X_{l})+\ldots+D_{n}(X_{k}),\,D_{n}(Y_{1})+\ldots+D_{n}(Y_{l}))+O_{k,l}(1).$
@@
Note that $\sum\nolimits_{i=1}^{l}D_{n}(X_{i})$ and $\sum\nolimits_{j=1}^{l}D_{n}(Y_{j})$ takes values in the dyadic rationals. One has
$H\left(\sum\limits_{i=1}^{k}D_{n}(X_{i}),\,\sum\limits_{j=1}^{l}D_{n}(Y_{j})\right)$ $=H\left(D_{n}\Big{(}\sum\limits_{i=1}^{l}D_{n}(X_{i})\Big{)},\,D_{n}\Big{(}\sum\limits_{j=1}^{l}D_{n}(Y_{j})\Big{)}\right)$
$=H_{n}\left(\sum\limits_{i=1}^{l}D_{n}(X_{i}),\,\sum\limits_{j=1}^{l}D_{n}(Y_{j})\right).$
The conclusion then follows from Lemma 2.18 with the $O_{k,l}(1)$-Lipschitz function given by
$(x_{1},\ldots,x_{k},y_{1},\ldots,y_{l})\mapsto(x_{1}+\ldots+x_{k},\,y_{1}+\ldots+y_{l}).$
For independent discrete random variables $X$ and $Y$, one has
$H(X+Y)\geq H(X+Y|Y)=H(X|Y)=H(X).$
Applying Lemma 2.19, for any independent random variables $X,Y$ and any $n\in\mathbb{N}$,
$H_{n}(X+Y)=H(D_{n}(X)+D_{n}(Y))+O(1)\geq H_{n}(X)+O(1).$
As a consequence,
$\max\{H_{n}(X),H_{n}(Y)\}\leq H_{n}(X+Y)+O(1)$ (13)
for all $n\in\mathbb{N}$, provided that $X$ and $Y$ are independent. In the rest of this paper, we may apply Lemma 2.19 without mention it.
In [16], Ruzsa studied some entropy inequalities involving sums of random variables. In the following, more Plünnecke-Ruzsa type inequalities for discretized entropy are presented.
@@
Let $X,Y$ and $Z$ be independent real random variables. Then, for all $n\in\mathbb{N}$,
@@
$H_{n}(X+Y+Z)+H_{n}(X,Y,Z)\leq H_{n}(X+Z,Y)+H_{n}(X,Y+Z)+O(1).$
Combining Lemma 2.4(iv) and the independence, the first inequality follows. Next, it follows from (13) that
$H_{n}(X+Y)\leq H_{n}(X+Y+Z)+O(1)\leq H_{n}(X+Z)+H_{n}(Y+Z)-H_{n}(Z)+O(1).$
Another application of Lemma 2.17 gives
@@
Combining independence, the second inequality involving XYX-Y then follows. ∎
@@
Let $X$ and $Y$ be independent real random variables and $k,l\in\mathbb{Q}$. There exist a non-negative integers $d=d(k,l)$ such that
$H_{n}(kX+lY)\leq dH_{n}(X+Y)-(d-1)(H_{n}(X)-H_{n}(Y))+O_{k,l}(1)$
for all $n\in\mathbb{N}$. Here by $kX$ we mean the product of $k$ and the random variable $X$.
@@
We first prove that
$H_{n}(kX+lY)\leq(d_{1}+d_{2}+1)H_{n}(X+Y)$
$\qquad-d_{1}H_{n}(X)-d_{2}H_{n}(Y)+(d_{1}+d_{2}+1)I_{n}(X,Y)+O_{k,l}(1)$ (14)
for some non-negative integers $d_{1}=d_{1}(k,l)$ and $d_{2}=d_{2}(k,l)$.
@@
by (13). Similar conclusion holds when $k=0$. In the following, we assume that $k,l\neq 0$.
@@
Indeed, when $k=l=1$, one can deduce (15) from (13). When $k=1$ and l=1l=-1, we take $X,Y,Z$ to be independent random variables such that $X$ and $Z$ are identically distributed. Then Corollary 2.20 gives
@@
Then (15) follows. Since $H_{n}(W)=H_{n}(-W)$ for any random variable $W$, the cases that k=1k=-1 and $l=\pm 1$ are also satisfied. Since $H_{n}(W)=H_{n}(-W)$ for any random variable $W$, the cases that k=1k=-1 and $l=\pm 1$ are also satisfied.
@@
where the random variable $Z$ is taken such that $X,Y,Z$ are independent and $X$, $Z$ are identically distributed. Combining (13), we obtain that
@@
Combining the fact
$H_{n}(X+Z)\leq H_{n}(X+Y)+H_{n}(Z+Y)-H_{n}(Y)+O(1)$
and the inductive hypothesis on $k$, one sees that the conclusion holds with some $d_{1}(k,l)$ and $d_{2}(k,l)$ satisfying
@@
By symmetry, we also have the induction on $l$, and the proof of (14) is completed for the situation that $k,l\in\mathbb{Z}^{\ast}$.
@@
for some $d_{1}=d_{1}(k^{\prime},l^{\prime})$ and $d_{2}=d_{2}(k^{\prime},l^{\prime})$. Then (14) follows.
Next, in view of (15), we can add multiple scalars of $H_{n}(X+Y)-H_{n}(X)$ and $H_{n}(X+Y)-H_{n}(Y)$ on the right-hand side of (14) to get
$H_{n}(kX+lY)\leq(2d_{1}+2d_{2}+1)H_{n}(X+Y)-(2d_{1}+2d_{2})\frac{H_{n}(X)+H_{n}(Y)}{2}+O_{k,l}(1).$
The lemma then follows by taking $d=2d_{1}+2d_{2}+1$. ∎
@@
for any $n\in\mathbb{N}$ and any i.i.d. random variables $X$, $Y$ with bounded range. However, when $k,l$ are $\mathbb{Q}$-linearly independent, such statement may fail. In Example 7.8, we would show the following: For any $K>0$, there exists some $N\in\mathbb{N}$ such that: for any $n>N$, there is a random variable $X$ with range contained in $[0,1]$ satisfying
@@ 2.4 Frostman measures
Now we recall the definitions of Frostman measures and show some of their properties.
Definition. Let $s\geq 0$ and $C>0$. We say that a Borel measure $\mu$ on $\mathbb{R}^{m}$ is $(s,C)$-Frostman if for any $x\in\mathbb{R}^{m}$ and $r>0$
@@
Definition.Let $s,s_{1},s_{2}\geq 0$ and $C,C_{1},C_{2}>0$.
(i) We say that the random variable $(X,Y)$ is jointly $(s_{1},s_{2},C)$-Frostman, if
$P(X\in B(x,r_{1}),\,Y\in B(y,r_{2}))\leq Cr_{1}^{s_{1}}r_{2}^{s_{2}}$
for any ball $B(x,r_{1})$ and $B(y,r_{2})$.
(ii) We say that the random variable $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, if $X$ and $Y$ are $(s_{1},C_{1})$-Frostman and $(s_{2},C_{2})$-Frostman, respectively.
(iii) We say a random variable $X$ is $(s,C)$-Frostman given $Y$, if
$P(X\in B(x,r),\,Y\in E)\leq Cr^{s}P(Y\in E)$
for any ball $B(x,r)$ and Borel set $E$. Moreover, We say that the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, if $X$ is $(s_{1},C_{1})$-Frostman given $Y$, and $Y$ is $(s_{2},C_{2})$-Frostman given $X$.
The following properties follows directly from the definitions, whose proofs are left as an exercise.
Lemma 2.22.
Let $s_{1},s_{2}\geq 0$ and $C,C_{1},C_{2},R>0$.
(i) Suppose that $X$ and $Y$ are independent random variables such that $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman. Then $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman.
(ii) Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman. Then it is both marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, and jointly $(s_{1},s_{2},C_{1}C_{2})$-Frostman.
(iii) Suppose that $(X,Y)$ is jointly $(s_{1},s_{2},C)$-Frostman. Then it is also $(s_{1}+s_{2},C)$-Frostman.
(iv) Suppose that $(X,Y)$ is a jointly $(s_{1},s_{2},C)$-Frostman random variable whose range is contained in $B(0,R)$, where $R>0$. Then $(X,Y)$ is marginally $(s_{1},CR^{s_{2}};s_{2},CR^{s_{1}})$-Frostman.
@@
Let $(X,Y)$ be a bivariate real random variable that is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, where $0\leq s_{1},s_{2}\leq 1$ and $C_{1},C_{2}>0$. Then the following statements hold for all $a,b,c,d\in\mathbb{R}$.
(i) The random variable $aX+bY$ is $(s_{1},C_{1}^{\prime})$-Frostman when $a\neq 0$, and is $(s_{2},C_{2}^{\prime})$-Frostman when $b\neq 0$. Here $C_{1}^{\prime},C_{2}^{\prime}$ are positive numbers satisfying $\log C_{i}^{\prime}=\log C_{i}+O_{a,b}(1)$ $(i=1,2)$.
(ii) Suppose that $ad-bc\neq 0$. Then (aX+bY,cX+dY)(aX+bY,\,cX+dY) is jointly $\big{(}s_{1},s_{2},C^{\prime\prime}\big{)}$-Frostman when $ad\neq 0$, and is jointly $\big{(}s_{2},s_{1},C^{\prime\prime}\big{)}$-Frostman when $bc\neq 0$. Here $C^{\prime\prime}$ is a positive number satisfying $\log C^{\prime\prime}=\log C_{1}C_{2}+O_{a,b,c,d}(1)$.